# Spend Wisely: Maximizing Post-Training Gains in Iterative Synthetic Data Boostrapping

### v1

@@ Abstract
Modern foundation models often undergo iterative “bootstrapping” in their post-training phase: a model generates synthetic data, an external verifier filters out low-quality samples, and the high-quality subset is used for further fine-tuning. Over multiple iterations, the model’s performance improves—raising a crucial question: how should the total budget on generation and training be allocated across iterations to maximize final performance? In this work, we develop a theoretical framework to analyze budget allocation strategies. Specifically, we show that constant policies fail to converge with high probability, while increasing policies—particularly exponential growth policies—exhibit significant theoretical advantages. Experiments on image denoising with diffusion probabilistic models and math reasoning with large language models show that both exponential and polynomial growth policies consistently outperform constant policies, with exponential policies often providing more stable performance.
Machine Learning, ICML
@@ 1 Introduction
A growing focus in the community is on post-training regimes, where models first acquire general knowledge during pretraining but require further refinement to excel in complex problem-solving tasks, such as agentic behavior (Shao et al., 2023), math reasoning (Poesia et al., 2024), and advanced code generation (Guo et al., 2024a). In these settings, the availability of high-quality human-annotated data often becomes a significant bottleneck due to the demand for expert input and substantial resources.
To address this limitation, synthetic data have emerged as a promising alternative, offering scalability and cost-effectiveness. Despite concerns about potential risks of model collapse (Shumailov et al., 2024; Dohmatob et al., 2024c), synthetic data is typically selected and verified before use in post-training (Feng et al., 2024; Setlur et al., 2024), ensuring its quality. A common paradigm for leveraging synthetic data employs an iterative bootstrapping process (Zelikman et al., 2022; Trung et al., 2024): the model generates synthetic data, rewards or verifiers are used to filter and select high-quality data, and the model is then fine-tuned on selected data. This process is repeated iteratively to fully improve performance. An illustration of this approach is provided in Figure 1.
However, for practitioners implementing this approach, an important question arises: how should a fixed computational budget be allocated to decide the amount of synthetic data to generate and select at each iteration to maximize performance?
@@
In this paper, we establish foundational principles for crafting optimal strategies for synthetic data generation across iterations. To the best of our knowledge, this is the first attempt to address this problem. We begin with a theoretical analysis of policies that control the amount of selected synthetic data used in each iteration, as illustrated in Figure 1. In a simplified setting with Gaussian data and exponential reward functions, we could identify the optimal policy. Contrary to common strategies that maintain a fixed number of synthetic data across iterations (constant policies), our analysis shows that the optimal strategy requires exponentially increasing the amount of data at each iteration.
@@
Building on these theoretical insights, we validate our findings experimentally through three settings: a toy example aligned with the previous simplified Gaussian data, an image-denoising task using Diffusion Probabilistic Models (DPM), and a math-reasoning task with Large Language Models (LLM). Across all experiments, both exponential and linear (polynomial) growth policies outperform constant policies. Moreover, the best exponential policies match or exceed the performance of linear policies in every case, confirming their theoretical robustness.
@@
  • We formulate the problem of optimizing post-training gains under a limited budget for iterative learning in Section 3.1.
  • In a solvable Gaussian setting, we prove the optimality of an exponential growth policy in Section 3.2 and validate it through numerical simulations in Section 4.1.
  • In a more general setting, we demonstrate in Section 3.3 that constant policies fail to converge to the optimal reward with high probability, while increasing policies ensure convergence. Among these, the exponential growth policy achieves an exponential convergence rate and outperforms polynomial growth policies in the worst case.
  • Experiments on image denoising (diffusion probabilistic model) and math reasoning (large language model) in Section 4 confirm that exponential and polynomial growth policies outperform constant policies, with exponential policies exhibiting greater stability.
@@ 2 Related Work
We review the relevant literature on synthetic data, with a focus on iterative bootstrapping and model collapse.
Synthetic Data and Iterative Boostrapping. The iterative bootstrapping framework, illustrated in Figure 1, is widely used for training foundational models in both text and vision domains. In this approach, synthetic data are generated and then iteratively refined or filtered to produce new training sets that more closely align with the target objectives. For large language models (LLMs), this approach has been applied to instruction following (Touvron et al., 2023) and alignment tasks (Dong et al., 2023), wherein reward models filter synthetic data to retain only high-quality samples. Recently, it has been particularly effective in improving reasoning abilities, as first demonstrated in (Zelikman et al., 2022). Subsequent works (Lehnert et al., 2024; Su et al., 2024; Trung et al., 2024; Guan et al., 2025; Singh et al., 2025) have extended this approach to generate diverse reasoning paths with search algorithms. Beyond text, iterative bootstrapping has also been utilized to generate training data for image segmentation (Kirillov et al., 2023), create image-text pairs (Fan et al., 2023, 2024), and synthesize images to address distribution shifts (Hemmat et al., 2024; Azizi et al., 2023). However, in these empirical works, a principled method for controlling generation in each iteration is still missing.
To our knowledge, Ferbach et al. (2024) provides the only theoretical analysis of this paradigm, proving convergence to the optimal policy under the infinite sample assumption. Crucially, their framework fails to characterize finite sample requirements, offering little actionable guidance to practitioners implementing this framework. We address this gap by establishing optimal computational budget allocation strategies across iterations for maximal performance gains.
Model Collapse. A series of papers has shown that incorporating synthetic data into the training corpus can lead to performance degradation, either in a single training instance or iteratively over time (Shumailov et al., 2023; Alemohammad et al., 2024; Dohmatob et al., 2024c, a, b). Feng et al. (2024) has explored the use of data selection to overcome model collapse by employing synthetic data in a single generation, combined with verification. In our work, we also leverage verification but focus on addressing the question of how to optimally generate synthetic data during learning.
3 Theoretical Results
In this section, we formalize the iterative learning setting and analyze optimal budget allocation strategies. We begin with the basic notation in Section 3.1 and a simple Gaussian case showing that exponential growth is optimal in Section 3.2. We then present our main theoretical results in Section 3.3, demonstrating that constant policies fail to converge with high probability, whereas there exists an exponential growth policy that achieves an exponential convergence rate and outperforms polynomial alternatives in the worst case.
@@ 3.1 Problem Formulation
We begin by formalizing the setup as illustrated in Figure 1. Our goal is to train a parameterized generative model $f(\cdot;\theta)$, such as an LLM or a diffusion probabilistic model (DPM). This model generates data $x\sim\mathcal{P}_{\theta}$, where $\mathcal{P}_{\theta}$ is the distribution induced by $f(\cdot;\theta)$. A reward model $\mathcal{R}$ evaluates the quality of the generated data, serving as a measure of the model’s performance. We aim to train the generative model to maximize the expected reward
@@
The iterative training process leverages the generative nature of the model to perform self-generation and refinement based on reward signals. Specifically, we start with an initial model $\theta^{(0)}$. At the $(t+1)$-th iteration, the model undergoes improvement via the following three steps:
  • Generation step: Generate $N_{t}$ data by $f(\cdot;\theta^{(t)})$.
  • Selection step: For each synthetic data point $x$, include it into dataset $D_{t}$ with probability $\mathcal{R}(x)$. Suppose $n_{t}$ data are selected such that $|D_{t}|=n_{t}$.
  • Updating step: Update $\theta^{(t)}$ to $\theta^{(t+1)}$ with $D_{t}$.
This iterative approach generates additional training data using synthetic samples, with the reward-guided selection process ensuring its quality. The selected samples are then used to further enhance the model’s generative performance (Feng et al., 2024). This process, carried out iteratively to maximize performance gains, is particularly effective in scenarios where high-quality specialized data is scarce—for example, in mathematical reasoning (Zelikman et al., 2022; Guan et al., 2025; Singh et al., 2025). The algorithm framework is formalized in Algorithm 1.
When practitioners use this algorithm, a key challenge is determining how to set $N_{t}$ and $n_{t}$ across iterations. This decision is particularly important as the generation cost of foundation models often exceeds the training cost. Given a limited budget, it is crucial to design the process in a way that maximizes post-training gains.
@@ 3.2 Warm-up
In Algorithm 1, we employ MLE as the updating algorithm at line 9. Specifically, the parameter $\theta^{(t+1)}$ updated at the $t$-th iteration is obtained by
@@
We aim to find the optimal iterative learning policy within this setup. We assume that the cost of sampling from the Gaussian distribution is negligible, while the training cost for MLE is proportional to the size of the dataset, $n_{t}$. The objective is to maximize the generator’s expected performance at iteration $T$, as measured by the reward. Given a fixed budget $C$, the optimal policy will be the solution to
$\begin{split}\max_{\{n_{t}\}_{t=0}^{T-1}}&\quad\mathbb{E}_{\theta^{(T)}}\left[r\left(\theta^{(T)}\right)\right],\\ \text{s.t.}&\quad\sum_{t=0}^{T-1}n_{t}\leq C.\end{split}$ (1)
@@
If the initial parameter satisfies $\theta^{(0)}\leq(1+\sigma^{2}/\kappa^{2})^{T}(\sigma^{2}+\kappa^{2})^{1/2},$ then the optimal iterative policy, i.e., the solution of the optimization problem in Equation 1, is given by
@@
The proof is deferred to Appendix A. Theorem 3.1 demonstrates that the optimal policy in this setting follows an exponential scheme. We further analyze this exponential scheme in the general analysis presented in the next section.
3.3 Main Theoretical Results
We now proceed to analyze the general setting. The assumptions and setup are outlined below, while detailed proofs for the subsequent theorems are provided in Section B.4.
@@ 3.3 Main Theoretical Results
We consider a general function class for the generative model, a general loss with regularity assumptions (assumption B.1, B.2), and a reward model with mild assumptions about its relationship with the loss (assumption B.3). The formal descriptions are provided in Appendix B. The warm-up example in the previous section serves as a special case that satisfies all these assumptions.
@@
In Algorithm 1, we use gradient descent with a learning rate $\eta>0$ as the updating algorithm at line 9. Specifically, the update for the $t$-th iteration writes
@@
where $l(x;\theta)$ is the loss function. For any policy $\pi$, let $\{\theta_{\pi}^{(t)}\}_{t}$ denote the iterative trajectory of the parameter.
@@
Similar to the warm-up example, we evaluate the generative model using the expected reward $r(\theta)$. The optimal performance is the supremum of the expected reward over the whole parameter space
@@
The first policy class we consider is
  • constant policy:
    $\pi_{\mathrm{const}}:n_{t}=n_{0},\quad t=0,1,\cdots.$
This is the most standard policy, where $n_{0}$ can either be set as the total number of prompts in the dataset (Zelikman et al., 2022), or as a multiple of the batch size in an online setting (Guo et al., 2024b). However, in the following theorem we show that constant policies fail to converge to the optimal reward with high probability.
@@
Under assumption B.3, there exists a constant $c>0$ such that, for any $T$ and any constant policy $\pi_{\mathrm{const}}$, with probability at least $1/4$,
@@
The theory establishes the non-convergence of constant policies, which stems from the persistent random noise in the gradient. Specifically, the noise term is proportional to $n_{t}^{-1}$ and remains non-decaying under any constant policy, resulting in suboptimal performance.
@@
Under assumption B.1 to assumption B.3, for any increasing policy $\pi$ and $\varepsilon>0$, there exists a sufficiently large $T$ such that, with probability greater than $1-\sum_{t=0}^{T-1}n_{t}^{-4}$,
@@
  • polynomial growth policy:
    $\pi_{\mathrm{poly}}:n_{t}=n_{0}(1+t)^{\alpha},\quad\alpha>0,\,t=0,1,\cdots,\text{and}$
  • exponential growth policy:
    $\pi_{\mathrm{exp}}:n_{t}=n_{0}(1+u)^{t},\quad u>0,\,t=0,1,\cdots.$
For the exponential policy, we identify a specific policy that achieves an exponential convergence rate as follows.
@@
Under assumptions B.1, B.2 and B.3, there exists an exponential growth policy $\pi_{\mathrm{exp}}^{*}$ such that with probability at least $1-\sum_{t=0}^{T-1}n_{t}^{-4}$,
@@
where $\zeta>0$ is a constant defined in the appendix.
We now compare the exponential policy with the polynomial policy. For any policy $\pi$, we introduce
@@
denoting the minimum number of iterations needed for the policy $\pi$ to achieve a reward that is within $\varepsilon$ of the optimal $r^{*}$. $C(\pi,T^{*}(\pi,\varepsilon))$ is then the minimum cost to attain the target performance for policy $\pi$. The following theorem compares the total cost incurred by the exponential growth policy and polynomial policies in a worst-case scenario.
@@
with probability at least $1-\sum_{t=0}^{T^{*}(\pi^{*}_{\mathrm{exp}},\varepsilon)-1}n_{t}^{-4}$. Here, the supremum is taken over all feasible problem settings, including parameter spaces, loss functions, and reward functions under mild assumptions.
For the formal version, please refer to Theorem B.7. This theorem shows that, in a worst-case analysis, the exponential policy $\pi_{\mathrm{exp}}^{*}$ from Theorem B.6 incurs a total cost no greater than that of any polynomial growth or constant policy to reach a specified performance threshold, establishing its efficiency and robustness. This analysis theoretically highlights the exponential policy as the preferred choice for our iterative learning tasks.
@@
Constant policies can result in rewards that are consistently below optimal by a measurable margin. Increasing policies, including polynomial and exponential growth ones, can achieve optimal performance. In the worst case, there exists one exponential growth policy that can guarantee the same performance at equal or lower computational cost compared to any constant or polynomial growth policy.
@@
Experiment Update Alg. Generator Reward Dataset
Toy Example MLE Gaussian Exponential Reward Gaussian Distribution
Image Denoising Adam DPM PSNR MNIST (Deng, 2012)
Math Reasoning Adam LLM Accuracy GSM-Symbolic (Mirzadeh et al., 2024)

### v2

# Spend Wisely: Maximizing Post-Training Gains in Iterative Synthetic Data Bootstrapping

@@ Abstract
Modern foundation models often undergo iterative “bootstrapping” in their post-training phase: a model generates synthetic data, an external verifier filters out low-quality samples, and the high-quality subset is used for further fine-tuning. Over multiple iterations, the model performance improves, raising a crucial question: How should the total budget for generation and training be allocated across iterations to maximize final performance? In this work, we develop a theoretical framework for analyzing budget allocation strategies. Specifically, we show that constant policies fail to converge with high probability, while increasing policies—particularly exponential growth policies—exhibit significant theoretical advantages. Experiments on image denoising with diffusion probabilistic models and math reasoning with large language models show that both exponential and polynomial growth policies consistently outperform constant policies, with exponential policies often providing more stable performance.
@@ 1 Introduction
Supervised fine-tuning is a critical stage in training large language models (LLM). Pre-training supplies them with broad linguistic and factual knowledge, but task-specific skills—such as tool use (Schick et al., 2023), reasoning pattern (Gandhi et al., 2025), and agentic behavior (Shao et al., 2023)—emerge when the model is further refined on carefully curated, supervised examples. Securing these high-quality human-annotated data, however, remains a major bottleneck, as it requires domain expertise and substantial resources.
To address this limitation, synthetic data have emerged as a promising alternative that offers scalability and cost-effectiveness. Despite concerns about potential risks of model collapse (Shumailov et al., 2024; Dohmatob et al., 2024b), synthetic data is typically selected and verified before use in post-training (Feng et al., 2024; Setlur et al., 2024), ensuring its quality. A common paradigm for leveraging synthetic data employs an iterative bootstrapping process (Zelikman et al., 2022; Trung et al., 2024): the model generates synthetic data, rewards or verifiers are used to filter and select high-quality data, and the model is then fine-tuned on selected data. This process is repeated iteratively to fully improve performance. An illustration of this approach is provided in Figure˜1.
However, for practitioners who implement this approach, an important question arises: How should a fixed computational budget be allocated to decide the amount of synthetic data to generate and select at each iteration to maximize performance?
@@
In this paper, we establish some foundational principles for crafting optimal strategies for synthetic data generation across iterations. To the best of our knowledge, this is the first attempt to address this problem. We begin with a theoretical analysis of policies that control the amount of selected synthetic data used in each iteration, as illustrated in Figure˜1. In a simplified setting with Gaussian data and exponential reward functions, we could identify the optimal policy. Contrary to common strategies that maintain a fixed number of synthetic data across iterations (constant policies), our analysis shows that the optimal strategy requires exponentially increasing the amount of data at each iteration.
@@
Building on these theoretical insights, we validate our findings with two experiments: an image-denoising task using diffusion probabilistic models (DPMs), and a math-reasoning task with large language models (LLMs). Across these experiments, exponential and linear (polynomial) growth policies outperform constant policies. Moreover, the best exponential policies match or exceed the performance of linear policies in every case, confirming their theoretical robustness.
@@
  • We formulate the problem of optimizing post-training gains under a limited budget for iterative learning in Section˜3.1.
  • In a solvable Gaussian setting, we prove the optimality of an exponential growth policy in Section˜3.2 and validate it through numerical simulations.
  • In a more general setting, we demonstrate in Section˜4 that constant policies fail to converge to the optimal reward with high probability, while increasing policies ensure convergence. Among these, the exponential growth policy achieves an exponential convergence rate and outperforms polynomial growth policies in the worst case.
  • Experiments on image denoising (diffusion probabilistic model) and math reasoning (large language model) in Section˜5 confirm that exponential and polynomial growth policies outperform constant policies, with exponential policies exhibiting greater stability.
@@ 2 Related Work
Synthetic Data and Iterative Bootstrapping. The iterative bootstrapping framework, illustrated in Figure˜1, is widely used for training foundational models in both text and vision domains. In this approach, synthetic data are generated and then iteratively refined or filtered to produce new training sets that more closely align with the target objectives. For large language models (LLMs) and vision language models (VLMs), this approach has been applied to instruction following (Touvron et al., 2023) and alignment tasks (Dong et al., 2023), and vision question answering (Yang & Dong, 2025), wherein reward models filter synthetic data to retain only high-quality samples. Recently, it has been particularly effective in improving reasoning abilities, as first demonstrated in Zelikman et al. (2022). Subsequent works (Lehnert et al., 2024; Su et al., 2024; Trung et al., 2024; Guan et al., 2025; Singh et al., 2025) have extended this approach to generate diverse reasoning paths with search algorithms. Beyond text, iterative bootstrapping has also been utilized to generate training data for image segmentation (Kirillov et al., 2023), create image-text pairs (Fan et al., 2023, 2024), and synthesize images to address distribution shifts (Hemmat et al., 2024; Azizi et al., 2023). However, in these empirical works, a principled method for controlling generation in each iteration is still missing.
To our knowledge, Ferbach et al. (2024) provides the only theoretical analysis of this paradigm, proving convergence to the optimal policy under the infinite-sample assumption. Crucially, their framework fails to characterize finite sample requirements, offering little actionable guidance to practitioners implementing this framework. We address this gap by establishing optimal computational budget allocation strategies across iterations to achieve maximum performance gains.
Model Collapse. A series of papers has shown that the incorporation of synthetic data into the training corpus can lead to performance degradation, either in a single training instance or iteratively over time (Shumailov et al., 2023; Alemohammad et al., 2024; Dohmatob et al., 2024b, a, 2025). Feng et al. (2024) has explored the use of data selection to overcome model collapse by employing synthetic data in a single generation, combined with verification. In our work, we also leverage verification, but focus on addressing the question of how to optimally generate synthetic data during learning.
3 Problem Formulation and Preliminary Analysis
In this section, we formalize the iterative learning setting in Section˜3.1 and then analyze optimal budget allocation strategies with a simple Gaussian case in Section˜3.2.
@@ 3.1 Problem Formulation
The setup is illustrated in Figure˜1. Our goal is to train a parameterized generative model $f(\cdot;\theta)$, for example, given by a DPM or an LLM, that generates data $x\sim\mathcal{P}_{\theta}$. A reward model $\mathcal{R}$ evaluates the quality of the generated data, serving as a measure of the model performance. We aim to train the generative model to maximize the expected reward
@@
The iterative training process leverages the generative nature of the model to perform self-generation and refinement based on reward signals. Specifically, we start with an initial model $\theta^{(0)}$. At the $(t+1)$-th iteration, the model undergoes improvement through the following three steps:
Generation step: Generate $N_{t}$ data by $f(\cdot;\theta^{(t)})$. Selection step: For each synthetic data point $x$, include it into dataset $D_{t}$ with probability $\mathcal{R}(x)$. Suppose $n_{t}$ of the $N_{t}$ data are selected to form the dataset $D_{t}$, formally, $|D_{t}|=n_{t}$. Updating step: Update $\theta^{(t)}$ to $\theta^{(t+1)}$ with the dataset $D_{t}$.
This iterative approach generates additional training data using synthetic samples, with the reward-guided selection process ensuring its quality. The selected samples are then used to further enhance the model’s generative performance (Feng et al., 2024). This process, carried out iteratively to maximize performance gains, is particularly effective in scenarios where high-quality specialized data is scarce, for example, in mathematical reasoning (Zelikman et al., 2022; Guan et al., 2025; Singh et al., 2025). The algorithm framework is formalized in Algorithm 1. In the selection step, the data may be selected with noise. We present a simple form here for ease of understanding, while our results extend to the noisy case in Section˜B.5.
When practitioners use this algorithm, a key challenge is determining how to set $N_{t}$ and $n_{t}$ across iterations. This decision is particularly important, as the generation process of foundation models unually involves multiple forward propagations, and consequently, the generation cost often exceeds the training cost. Given a limited budget, it is crucial to design the process in a way that maximizes post-training gains.
@@ 3.2 Gaussian Case Study
In Algorithm˜1, we employ MLE as the updating algorithm at line 9. Specifically, the parameter $\theta^{(t+1)}$ updated in the $t$-th iteration is obtained by
@@
We aim to find the optimal iterative learning policy within this setup. We assume that the cost of sampling the Gaussian distribution is negligible, while the training cost for MLE is proportional to the size of the dataset, $n_{t}$. The objective is to maximize the expected performance of the generator in iteration $T$, as measured by the reward. Given a fixed budget $C$, the optimal policy will be the solution to
$\max_{\{n_{t}\}_{t=0}^{T-1}}\mathbb{E}_{\theta^{(T)}}\left[r\left(\theta^{(T)}\right)\right]\quad\text{s.t.}\quad\sum_{t=0}^{T-1}n_{t}\leq C.$ (1)
@@
If the initial parameter satisfies $\theta^{(0)}\leq(1+\sigma^{2}/\kappa^{2})^{T}(\sigma^{2}+\kappa^{2})^{1/2},$ then the optimal iterative policy, i.e., the solution of the optimization problem in Equation˜1, is given by
@@
The proof is deferred to Appendix˜A and extends naturally to the high-dimensional case, as detailed in Section˜E.1. Theorem˜3.1 demonstrates that the optimal policy in this setting follows an exponential scheme. The optimal strategy follows this rationale: In an iterative optimization algorithm, the contribution of an early update to the overall error decays exponentially with the total number of iterations $T$. As a result, achieving a low final error increasingly depends on the accuracy of later updates. To keep those updates precise enough, each iteration must therefore draw on an exponentially larger number of samples. We will further analyze this exponential scheme in the general analysis in Section˜4.
Experiments.
We include an empirical experiment to confirm Theorem˜3.1 with data in two-dimensional space. The generator is set to $\mathcal{N}(\bm{\theta},\bm{I}_{2})$. The reward function is $\mathcal{R}(\bm{x})=\exp\left(-\lVert\bm{x}\rVert^{2}/(2\kappa^{2})\right)$, which favors data concentrated near the origin. $\kappa$ controls the flatness of the reward. Full details are deferred to Section˜E.2. Figure˜2 plots the gap to the optimal reward as a function of the cumulative cost, defined as the total computational cost incurred up to that point. The results demonstrate that the exponential policy consistently approaches the optimal reward faster than the other schemes. While the linear policy converges slightly more slowly, the constant policy, as shown in the left figure, clearly fails to converge. These findings validate our theoretical proof. Additional results with various parameters are left in Section˜F.1.
Figure 2: Empirical results of the toy example. We compare the exponential, constant, and linear policy, and show the gap to the optimal expected reward as a function of the computational cost ($\sum_{t}n_{t}$). All the results are averaged over 1,000 runs.
4 Main Theoretical Results
We now proceed to analyze the general setting. We will show that constant policies fail to converge with high probability, whereas there exists an exponential growth policy that achieves an exponential convergence rate and outperforms polynomial alternatives in the worst case. The assumptions and setup are outlined below, while detailed proofs for the subsequent theorems are provided in Section˜B.4.
@@ 4 Main Theoretical Results
We consider a general function class for the generative model, a general loss with regularity assumptions (Assumptions˜B.1 and B.2), and a reward model with mild assumptions about its relationship with the loss (Assumption˜B.3). The formal descriptions are provided in Appendix˜B. The warm-up example in the previous section serves as a special case that satisfies all these assumptions. Our results also naturally extend to noisy rewards, as discussed in Section˜B.5.
@@
In Algorithm˜1, we use gradient descent with a learning rate $\eta>0$ as the update algorithm at line 9. Specifically, the update for the $t$-th iteration writes
@@
where $l(x;\theta)$ is the loss function. For any policy $\pi$, let $\left\{\theta_{\pi}^{(t)}\right\}_{t}$ denote the iterative trajectory of the parameter.
@@
Similarly to the warm-up example, we evaluate the generative model using the expected reward $r(\theta)$. The optimal performance is the supremum of the expected reward over the whole parameter space
@@
The first policy class that we consider is
constant policy: $\pi_{\mathrm{const}}:n_{t}=n_{0},\quad t=0,1,\cdots.$
This is the most standard policy, where $n_{0}$ can be set either as the total number of prompts in the dataset (Zelikman et al., 2022), or as a multiple of the batch size in an online setting (Guo et al., 2024). However, in the following theorem we show that constant policies fail to converge to the optimal reward with high probability.
@@
Under Assumption˜B.3, there exists a constant $c>0$ such that, for any $T$ and any constant policy $\pi_{\mathrm{const}}$, with probability at least $1/4$,
@@
The theory establishes the non-convergence of constant policies, which stems from persistent random noise in the gradient. Specifically, the noise term is proportional to $n_{t}^{-1}$ and remains non-decaying under any constant policy, resulting in suboptimal performance.
@@
Under Assumption˜B.1 to Assumption˜B.3, for any increasing policy $\pi$ and $\varepsilon>0$, there exists a sufficiently large $T$ such that, with probability greater than $\left(1-\sum_{t=0}^{T-1}n_{t}^{-4}\right)$,
@@
polynomial growth policy: $\pi_{\mathrm{poly}}:n_{t}=n_{0}(1+t)^{\alpha},\quad\alpha>0,\,t=0,1,\cdots;$ exponential growth policy: $\pi_{\mathrm{exp}}:n_{t}=n_{0}(1+u)^{t},\quad u>0,\,t=0,1,\cdots.$
For the exponential policy, we can identify a specific policy that achieves an exponential convergence rate as follows.
@@
Under Assumptions˜B.1, B.2 and B.3, there exists an exponential growth policy $\pi_{\mathrm{exp}}^{*}$ such that with probability at least $\left(1-\sum_{t=0}^{T-1}n_{t}^{-4}\right)$,
@@
where $\zeta>0$ is a constant defined in Theorem˜B.6.
This theorem also applies to noisy rewards, as discussed in Section˜B.5.
Notice that the above three theorems analyze the convergence of the respective policies (constant, polynomial, and exponential) as the number of iterations $T$ grows without imposing any limit on the total computational budget. We now introduce the budget constraint and compare the exponential policy with the polynomial policy. Specifically, for any policy $\pi$, we introduce
@@
denoting the minimum number of iterations needed for the policy $\pi$ to achieve a reward that is within $\varepsilon$ of the optimal $r^{*}$. The cost $C(\pi,T^{*}(\pi,\varepsilon))$ is then the minimum cost to attain the target performance of the policy $\pi$. The following theorem compares the total cost incurred by the exponential growth policy and polynomial policies in a worst-case scenario.
@@
with probability at least $\left(1-\sum_{t=0}^{T^{*}(\pi^{*}_{\mathrm{exp}},\varepsilon)-1}n_{t}^{-4}\right)$. Here, the supremum is taken over all feasible problem settings, including parameter spaces, loss functions, and reward functions under mild assumptions.
For the formal version, please refer to Theorem˜B.7. This theorem shows that, in a worst-case analysis, the exponential policy $\pi_{\mathrm{exp}}^{*}$ from Theorem˜B.6 incurs a total cost no greater than that of any polynomial growth or constant policy to reach a specified performance threshold, establishing its efficiency and robustness. This analysis theoretically highlights the exponential policy as the preferred choice for our iterative learning tasks.
@@
Constant policies lead to rewards that are consistently below optimal by a constant margin. Increasing policies such as polynomial and exponential ones can achieve optimal performance. In the worst case, there exists an exponential growth policy ensuring the same performance at equal or lower computational cost compared to any constant or polynomial growth policy.
@@
Experiment Setup Type Generator Reward Dataset
Image Denoising Theoretical Validation DPM PSNR MNIST (Deng, 2012)
Math Reasoning Practical Scenario LLM Accuracy GSM-Symbolic (Mirzadeh et al., 2025)